### v4

# TAS: A Transit-Aware Strategy for Embodied Navigation with Non-Stationary Targets

Abstract
Embodied navigation methods commonly operate in static environments with stationary targets. In this work, we present a new algorithm for navigation in dynamic scenarios with non-stationary targets. Our novel Transit-Aware Strategy (TAS) enriches embodied navigation policies with object path information. TAS improves performance in non-stationary environments by rewarding agents for synchronizing their routes with target routes. To evaluate TAS, we further introduce Dynamic Object Maps (DOMs), a dynamic variant of node-attributed topological graphs with structured object transitions. DOMs are inspired by human habits to simulate realistic object routes on a graph. Our experiments show that on average, TAS improves agent Success Rate (SR) by 21.1 in non-stationary environments, while also generalizing better from static environments by 44.5% when measured by Relative Change in Success (RCS). We qualitatively investigate TAS-agent performance on DOMs and draw various inferences to help better model generalist navigation policies. To the best of our knowledge, ours is the first work that quantifies the adaptability of embodied navigation methods in non-stationary environments. Code and data for our benchmark will be made publicly available.
1 Introduction
Embodied navigation involves an agent moving in unseen physical or virtual environments to carry out human-guided tasks such as locating objects [4, 36, 10, 33], following instructions [19, 13, 1, 31], or reaching a point in space [49, 45]. State-of-the-art methods in this field use Reinforcement Learning (RL)  [47, 45, 20] and Large-Language Model (LLM) based schemes [39, 10] to achieve superior task performance, and these policies also extend to multi-object finding [9, 44]. Topological graphs (TGs) are a common facet of such techniques and are used to store environmental details for decision making. Embodied instruction following tasks such as REVERIE [31]and R2R [1, 13] utilize TGs extracted from MP3D [8] for navigation. Similarly, tasks like ObjectNav, PointNav, and ImageNav also employ TGs [20, 46, 14, 48, 11, 22, 34] where nodes represent locations or landmarks, and edges define traversable paths. These graphs provide a structured way to plan and execute navigation policies.
A key feature of current embodied navigation paradigms [44, 9, 18, 12] is their reliance on static graphs, where objects within nodes remain stationary. This static assumption is limiting, as real-world environments are inherently dynamic, where users frequently move portable objects from place to place over long periods, such as shifting their phone or wallet between rooms throughout the day. Studies on robots in collaborative environments  [29, 2, 16] suggest that human object placements follow structured patterns shaped by routines and habits, introducing a degree of entropy in object locations over time. Despite the important applicability of non-stationary environments, existing navigation methods fail to model such temporal variability, limiting their applicability for real-world deployment.
Figure 1: Transit-Aware Strategy (TAS) We introduce TAS to tackle embodied navigation in non-stationary environments. In this figure, an embodied agent is tasked with finding a wallet that changes positions in the environment between 9:$0$ and 9:20. $P_{\text{wallet}}$ represents the transit route of the wallet. TAS makes agents “transit-aware” by modeling object transit information to augment the agent’s navigation policy for temporally-informed decision making.
In the ObjectNav task [18, 3], for instance, SotA navigation approaches all involve finding stationary targets using RL [47, 45, 3, 41] and LLM-based zero-shot methods [18, 39, 6]. RL policies for multi-object settings [44, 9] have also been proposed, but these are also based on static environments. In contrast, navigating to non-stationary targets that move as the agent moves necessitates the temporal modeling of object transit to guide the agent.
For example, OVRL-V2 [47] uses CLIP text embeddings and ViT-processed RGB images, while DD-PPO [45] similarly relies on GPS readings and RGB-D sensors; both use only static observations of the scene. Their reward functions similarly assume target stationarity, with OVRL-V2 and DD-PPO penalizing absolute distance changes to the goal. Using these approaches directly in non-stationary environments shows poor adaptability because the distance to a target fluctuates as the agent moves, producing a high-variance signal that hinders policy convergence.
For real-world deployment, these approaches must generalize, i.e., maintain performance in dynamic scenes with shifting targets. Prior embodied work describes generalizability in terms of measuring an agent’s capacity to adapt to novel settings, but this usually refers to improved performance in unseen environments or unique synthesized guidance language [50, 25, 23, 17]. Little to no work that defines generalizability in the context of adapting to dynamic settings with non-stationary targets.
Main Results. To address these issues, we present a Transit-Aware Strategy (TAS) that informs embodied navigation policies with temporal “awareness” on target positions (Fig. 1). Rather than naively chasing target objects, TAS augmented agents learn to intercept target routes for better convergence. To evaluate TAS, we introduce Dynamic Object Maps (DOMs), a dynamic variant of topological graphs incorporating structured object transitions. Figure 2 illustrates this concept. DOM node attributes change over time, i.e. target objects move from node to node at each timestep, while the edges remain fixed.A successful agent must synchronize its path with these object routines, which requires learning and anticipating object movement patterns during exploration. This ability reflects an agent’s generalization to non-stationary environments.
Figure 2: Dynamic Object Maps (DOMs): We introduce dynamism to topological graphs via portable targets following human-habit inspired routes. In this representative figure, the watch (in red) moves from node N1 at T=10:$0$ AM to node N5 at T=7:45 PM, while the wallet (in orange) moves from node N5 at T=10:15 AM to node N1 at T=7:45 PM. We benchmark navigation agents under different object transition scenarios and present a Transit-Aware Strategy (TAS) to improve their performance.
Our work seeks to redefine success as defined in traditional embodied navigation tasks, by emphasizing not just on where to navigate but also when to get there i.e. right place, right time. Our contributions are summarized as follows:
  • Transit-Aware Strategy (TAS): We propose TAS, a novel strategy to enhance the capabilities of embodied agents to find non-stationary targets. TAS equips an agent with object transit routes, enabling them to synchronize their paths with moving targets. We develop two TAS-enhanced agents TAS-RL and TAS-LLM, that modify existing RL [45] and LLM [18] policies with temporal transit cues.
  • Dynamic Object Maps (DOMs): We propose an algorithm to introduce structured object transitions on static topological graphs, transforming them into Dynamic Object Maps or DOMs. Inspired by human habits, we define three distinct object transition scenarios to govern the movement of portable objects in a scene, providing varying levels of entropy. DOMs apply to standard TGs from datasets like Matterport3D (MP3D) [8] and HM3D [32].
  • Generalizability Benchmark: We evaluate $6$ different navigation agents on a multi-object finding task across static and DOM environments. Our TAS-RL and TAS-LLM agents show an average improved Success Rate (SR of $22.9\%$ and generalize better by $49.8\%$ in terms of Relative Change in Success (RCS)[17] over non-TAS counterparts.
2 Related Works
Traditionally, navigation in dynamic environments has focused on low-level planning tasks, such as crowd avoidance [37, 27, 5] and socially aware navigation [7, 26], where agents continuously maneuver to avoid obstacles while minimizing trajectory length or time. In contrast, embodied navigation tasks employing high-level planning on graphs are predominantly studied in static environments with stationary targets, with limited research on handling dynamic targets  [13, 38]. In our work, we introduce dynamism on a graph-level, where agents must reason over long-term spatio-temporal changes of objects for navigation decisions. While recent simulators like Habitat 3.0 [30] include object rearrangement tasks, these primarily focus on agents modifying the environment and not finding non-stationary targets on a graph.
Planning in dynamic environments has been studied in the past [42, 52, 28], with recent approaches even utilizing LLMs in conjunction with multi-arm bandits [15]. These schemes usually propose memory augmentations with hierarchical planning procedures and frame the problem from an obstacle avoidance standpoint [42, 24, 51]. In contrast, our work considers the case where the objects themselves are non-stationary.
Rudra et. al. [36] define small portable objects around the house and propose a contextual bandit scheme that aims to learn the likelihood of finding an object at various waypoints. In their case, however, object locations are shuffled only after each episode, meaning it finally boils down to finding stationary targets at each timestep. In contrast, we tackle a truly dynamic case, where objects move even during the episode. This definition adds a layer of complexity as the embodied agent must now plan to navigate towards a shifting target, requiring an explicit modeling of the object’s transit.
Kurenkov et. al. [21] introduced dynamizing household environments and experimented with scene graph memory to predict object locations. They also performed target object-finding experiments, but the target was stationary as the agent moved. In our work, the objects move as the agent moves, and we deal with topological graphs, not scene graphs. Wang et. al in [43] used LLM-generated human activities to dynamize a topological graph of a household environment and performed ObjectNav experiments in these environments. In contrast, our work presents a transit-aware strategy to improve navigation performance in non-stationary environments.
3 Method: Finding Non-Stationary Targets
In this section, we first our approach to developing an environment with non-stationary targets with human-habit inspired transit scenarios. We then present our strategy to improve embodied navigation performance on these environments by integrating agents with transit-awareness.
3.1 Defining Transit Scenarios $\Lambda$:
Human habit formation via object placement can be attributed to cognitive offloading, where individuals rely on the environment to reduce memory demands [35, 40]. This reliance on the environment influences the spatial distribution of objects. Frequently used objects (like keys or phones) might move more flexibly across rooms (high entropy), while function-specific objects (like dumbbells or toothbrush) might be more constrained (low entropy). We use this structured transition behavior to define an object’s transition entropy, which we use to formulate $\Lambda$. We consider three object transit scenarios at decreasing levels of transit entropy, Random, Semi-Routine, and Fully-Routine. These are outlined in Table 1 below.
Scenario ($\mathbf{\Lambda}$) Fixed Rooms Fixed Paths Entropy
Static N/A None
Random High
Semi-Routine Medium
Fully-Routine Low
Table 1: Object Transit Scenarios $\mathbf{\Lambda}$: Portable targets transition in the graph under structured scenarios inspired by human habits. In the random case, the portable objects can move to any room at any time during each episode. In the routine cases, the rooms that the target objects can travel to are fixed (except during transit). The static scenario is a baseline with stationary items i.e., zero entropy.
Figure 3: Object Transition: Portable objects move around the scene at various timesteps in accordance to their natural rooms (Table I in Appendix) and transit scenarios (Table 1). Here, a mug is placed on a kitchen node $N_{3}$ at timestep $T=53$, but moves to a bedroom node $N_{5}$ at timestep $T=55$ via node $N_{4}$. If an agent reaches the kitchen after $T=53$ (or the bedroom before $T=55$), it would fail to see the mug. Multiple objects could also be at the same node (hat and mug in $N4$).
3.2 Dynamic Object Maps (DOMs):
We define Dynamic Object Maps (DOMs) as node-attributed dynamic graphs where edge values remain constant, but the node attributes (objects) evolve over time. DOMs extend attributed graphs with temporal object dynamics. Formally, let $G=(V,E)$ be a topological graph structure with nodes $V$ representing spatial locations (rooms and objects), and edges $E$ representing traversable paths between the nodes. Let $\mathcal{O}$ represent the total set of available portable objects. Then, at each timestep $t$, $v\in V$ is associated with a set of portable target objects $O_{v}(t)\subseteq\mathcal{O}$. While the graph edges $E$ remain fixed over time, the node attributes or object assignments $O_{v}(t)$ evolve via a time-dependent transit scenario $\Lambda$, to give us an object transition function $O$ as,
$O:V\times\Lambda\rightarrow 2^{\mathcal{O}}$ (1)
where $2^{\mathcal{O}}$ is a power set representing all possible portable object combinations. At each timestep $t$, $O(v,t)$ gives us the portable targets present in each node $v$ based on the transit scenario $\Lambda$.
Figure 3 illustrates this evolving graph setup. The initial graph contains portable objects placed according to Table I (See Appendix). We then generate object paths $\zeta$ by moving them in the graph at various timesteps for each transit scenario $\Lambda$. Algorithm I in the Appendix describes our approach to generating $\zeta$.
When generating the evolving graph, a pseudo-random seed $s$ helps differentiate the movement of portable objects per episode. Letting $e$ be the episode index,
[…]
Note that for the fully routine case, $\zeta$ remains constant for all episodes of the experiment, meaning that the object transition routes remain fixed. In the other two cases, $\zeta$ varies per episode according to the room and interval variability described.
3.3 Transit-Aware Strategy (TAS)
Many prior embodied navigation tasks work under the premise of “static and stationary”, where targets are assumed to stay at one location throughout the episode. This premise is also true with zero-shot LLM-based approaches [39, 18] As such, they are trained with a reward to penalize the agent for being too far from the target [45, 47]. This can be written as -
$\mathcal{R}_{\text{dist}}=-\alpha\parallel{v_{a}-v_{o}}\parallel.$
This standard formulation fails with DOMs, since $v_{o}$ changes during the episode. As a result, even though $R_{\text{dist}}$ is a dense reward, the reward signal fluctuates too much to enable meaningful learning. This fluctuation is reflected in the poor performance of DD-PPO [45] on DOMs and is discussed later in the results section.
To improve performance, agents need to be made “aware” of target objects in transit, and this involves integrating them with temporal knowledge of the scene. We present modifications to an RL agent and an LLM agent as follows:
TAS-RL Agent: To make an RL agent transit-aware, we make the following modifications to its setup:
Observations and Actions: Given a DOM $G=(V,E)$, we provide the agent with a set of paths $\mathcal{P}$ denoting potential object transits. $P_{o}\in\mathcal{P}$ is a list of timed vertices $v_{1},\dots v_{n}]$ denoting the position of object $o$ between timesteps $1\dots n$. Formally,
$\mathcal{P}=\left\{P_{o}\;\middle|\;o\in\mathcal{O}\right\},\quad\text{where}\quad P_{o}=\left\{(t_{i},v_{i})\right\}_{i=1}^{n},\;t_{i}\in T,\;v_{i}\in V.$
When $\mathcal{P}$ is not readily available, it can be initialized with known priors about human habits in the environment. In a practical example, if a human moves tends to move their wallet between the kitchen and bedroom between 99:$00$ and 99:2020, the wallet path could be approximated as $P_{wallet}=[v_{9:00},v_{9:01}\dots v_{9:20}]$, where $v_{t}$ indicates the wallet’s position a time $t$. The transit scenario $\Lambda$ categorizes $P_{o}$’s temporal structure.
We also provide the current timestep $t\in T$ and the current position of the agent $v_{c}$. Given this information, the agent chooses the next node to go to. We employ a maskable action space to ensure that the agent moves only to the neighboring nodes of $v_{c}$.
Reward Objective: As described earlier, a distance reward $R_{dist}$ provides a poor and highly variable training signal. We replace it with a sparse interception reward $\mathcal{R}_{I}$ that is triggered whenever the agent intersects with the target’s path $P_{o}$.
$\mathcal{R}_{I}=\begin{cases}r_{i},\>\text{if}\>v\in P_{o}\\ 0,\>\text{otherwise}.\end{cases}$
This reward incentivizes the agent to synchronize its path with that of the target, rather than greedily “chasing” a target to minimize distance. $\mathcal{R}_{I}$ along with a standard target finding reward $\mathcal{R}_{f}$ allows for more informed agent pathing for finding non-stationary targets.
TAS-LLM Agent: LLM-based navigation agents use commonsense reasoning about observations to predict navigable actions. In LGX [18] for instance, if the target is a ‘leaf-shaped bowl’, the agent uses observations of regular objects like ‘table’ and ‘sofa’, to determine a direction that is more likely to have the target (the ‘table’).
In DOMs, object transit scenarios $\Lambda$ similarly use commonsense locations defined in Table I (See Appendix). As such, passing previously seen regular objects to the LLM would help the agent make more transit-aware decisions. Since LLM agents are predominantly zero-shot and work directly during inference time with no training, we modify the system prompt to make them transit-aware.
At each timestep $ts_{i}$, we gather the action sequence $a_{0->ts_{i}}$, and a set of objects in the scene (may also include the target) $o_{t}$. In the subsequent timesteps, we pass this transit data $\mathcal{T}_{i}=[ts_{i},a_{0->ts_{i}},o_{t}]$ into the system prompt. A horizon is maintained to avoid token overflow, where the earliest appended observation is removed at the horizon.
For both TAS-RL and TAS-LLM, note that the Transit Scenario $\Lambda$ plays an important role in determining agent success. An agent finding an object in random transit $\Lambda_{\text{random}}$ would find it a lot harder than one in the routine transit cases $\Lambda_{\text{semi-routine}}$ or $\Lambda_{\text{routine}}$. For the former, even if an agent intersects with $\Lambda_{\text{random}}$ and follows the trajectory, it may never find the target is $\Lambda_{\text{random}}$ is spatio-temporally random.
4 Task: Multi-Object Finding
To evaluate the benefits of TAS, we focus on the task of finding multiple portable target objects over a fixed timespan. This setting adds realism to standard multi-object finding tasks such as Multi-ON [44] and GOAT [9] where targets are stationary. In our dynamic setting, we argue that a multi-object finding task is preferable over finding a single portable object for two reasons:-
  • Realistic Setting: In real-world homes, users frequently move multiple objects simultaneously, leading to dynamically changing object locations [2, 29]. Assistive agents must track and infer object movements over time to provide meaningful assistance when asked. This requires monitoring multiple portable objects (e.g., phones, watches) rather than treating each search as an isolated event.
  • Lifelong Navigation: A multi-object setup enables an agent to finetune its search policy in real-time, by using knowledge gathered from finding previous objects to better find the next. An agent deployed at homes will likely have to face similar conditions, with a user asking it to find multiple targets without ‘resetting’ its position each episode as the case with single-target tasks, much akin to lifelong navigation[46].
First, we mathematically formalize what navigating on a DOM entails. Let $\Lambda$ describe the transitory motions of objects. Then, let $O$ be the set of portable objects found over $T$ timesteps for an agent starting at node $r$. Then $S=[r,T,\Lambda]$ represents a set of our experimental variables for this task. We then define our navigation objective as “Finding the maximum possible number of portable objects $O$ while traversing a dynamic environment represented by SS.”.
For downstream applications like embodied agents personalizing to households, we believe finding multiple targets to be a good task to study dynamic adaptability, due to dense reward signals [9]. The setting is also realistic as households have multiple people moving objects around at the same time. We study the performance of different navigation policies with varying simulator conditions $S$.
In each episode, the agent is tasked with finding as many portable target objects as possible by making decisions on a dynamic DOM. Let $\tau$ be a finite trajectory representing a sequential list of visited nodes on the DOM, $G$, and let $\tau(t)$ be the $t$-th node in $\tau$. We can now define $O(\tau,t)$ as the set of portable objects found at the $t$-th node of $\tau$ at time $t$. Given that the objects move throughout the scene, the $t$-th node may have a different set of portable objects present at a different timestep. Given a trajectory $\tau$ of length $T$, we can formulate the total set of portable objects found along $\tau$ as $O_{\tau}=\cup_{t=1}^{T}O(\tau,t)$.
A navigation agent parameterized by policy $\pi$ explores the dynamic DOM by generating a trajectory $\tau_{\pi}$, where each timestep $t$ corresponds to a selected node: $\tau_{\pi}(t)\in V$. Specifically, $\tau_{\pi}(t)$ is the node chosen at time $t$ by policy $\pi$. We now can write the policy optimization problem as finding an optimal policy $\pi^{*}$ defined as $\pi^{*}=\operatorname*{arg\,max}_{\pi}|O_{\tau_{\pi}}|$. Consequently, $O_{\tau_{\pi^{*}}}$ is the maximum number of portable targets that can be found by taking the optimal trajectory.
5 Experiments and Results
We perform multi-object finding on the $4$ different $\Lambda$ scenarios defined in Table 1. DD-PPO [45] and LGX [18] are baseline RL and LLM agents and are TAS enhanced using the strategy described in Section 3.3 to produce TAS-RL and TAS-LLM agents. We also benchmark performance with Random and Oracle agents.
5.1 DOM Setup
For our experiments, we convert static topological graphs from the Matterport3D (MP3D) [8] dataset to obtain DOMs. Each node contains panoramic images representing the scene, and edges represent the distance between them. We determine $4$ hyperparameters that influence performance i.e., $H=[r,E,T,\Lambda]$, where $r\in G$ is a random starting node chosen from $G$, $E$ is is the number of episodes per trial, $T$ is the number of timesteps per episode and $\Lambda$ is the transit scenario. Thus, each agent is subjected to a total of $Stationary(1)+Transit(3)]\times r\times E\times T$ experiments. We define each trial as a set of episodes over a fixed time from a starting node $r$. The maximum possible number of trials for a scan is thus equal to the size of the graph $G$. In our subset of $10$ MP3D scans (Ref. Appendix for details), $|r|$ ranges from $[20,215]$. For each scan, we set $|r|=10$, $E=20$, and $T=30$ in our experiments.
Although different objects can have different transition scenarios, we set all objects in our experiments to have the same $\Lambda$. This allows us to better analyze the generalizability of a navigation scheme and avoids complexities involved with defining per-object $\Lambda$ associations.
Ground Truth Paths ($\Pi^{*}$): Given a starting node $r\in G$, for each episode $e\in E$, there exist multiple ground truth paths $\Pi^{*}$ over $T$ that the agent could take for collecting the most number of portable targets. We compute these paths by brute force, simulating all possible trajectories from $r$ over $T$ timesteps. $\Pi^{*}$ is used in our evaluation metrics.
5.2 Evaluation Metrics
Success Rate (SR): For each episode, we measure the object-finding performance of an agent by dividing the number of objects found in that episode by the maximum possible number of objects that could have been found with a ground-truth path $\pi^{*}\in\Pi^{*}$: $SR=|O_{\pi}|/|O_{\pi^{*}}|$. This metric is similar to the Progress metric in [44]. In our task, the ground-truth trajectory may not necessarily find all the objects in the given timeframe.
Trajectory Alignment (TA): To measure path efficiency, we define Trajectory Alignment as the average overlap between ground-truth trajectory $\pi^{*}$ and trajectory taken by the agent across episodes. We take the maximum across $\Pi^{*}$: $TA=\min_{\pi^{*}\in\Pi^{*}}\frac{1}{T}\sum_{t\in T}\mathbf{1}_{\pi(t)=\pi^{*}(t)}$, where $\mathbf{1}$ is an indicator function.
Relative Change in Success (RCS): We use this value taken from [17] to measure the generalizability from static to dynamic environments. It is the percent relative change between static and dynamic navigation performance. We measure RCS for various $\Lambda$ cases and report scores on SR and TA. Lower values indicate better generalizability, showing consistency in agent performance between static and dynamic environments.
Table 2: Navigation on DOMs: We evaluate 66 embodied navigation approaches on various DOMs to establish a benchmark. These values are used to compute RCS using the equation provided in [17]. An RCS value closer to $0\%$ indicates good generalizability to new conditions. Highlighted in green are the best RCS values for SR, while ones in violet are the best values for TA. Observe the improved RCS scores of TAS-RL and TAS-LLM agents, indicating better generalizability to non-stationary targets. Particularly note that the RL agent performs well in the static scenario, with a significant performance drop when subjected to DOMs. The LLM agent shows improved performance on DOMs, but we attribute this to multiple chance encounters with moving targets.
$\Lambda$ Random $\Lambda$ Semi-Routine $\Lambda$ Fully-Routine $\Lambda$ Average Policy Metric (%) Static DOM RCS Static DOM RCS Static DOM RCS Static DOM RCS Random SR $42.0$ $74.4$ +43.5+43.5 $42.0$ $54.7$ +23.2+23.2 $42.0$ $48.3$ +13.0+13.0 $42.0$ $59.1$ +29.0+29.0 TA $33.5$ $24.4$ 37.3-37.3 $33.5$ $36.6$ +8.4+8.4 $33.5$ $33.6$ +0.29 $33.5$ $31.2$ 6.8-6.8 Oracle SR $90.2$ $78.4$ 13.0-13.0 $90.2$ $75.1$ 16.8-16.8 $90.2$ $55.9$ 38.0-38.0 $90.2$ $69.8$ 22.6-22.6 TA $45.6$ $54.4$ +16.2+16.2 $45.6$ $50.6$ +9.8+9.8 $45.6$ $32.7$ 28.3-28.3 $45.6$ $45.9$ +0.6 RL [45] SR $61.8$ $33.4$ 45.9-45.9 $61.8$ $21.8$ 64.7-64.7 $61.8$ $12.2$ 80.1-80.1 $61.8$ $22.4$ 63.7-63.7 TA $59.8$ $36.5$ 38.9-38.9 $59.8$ $29.8$ 50.1-50.1 $59.8$ $24.1$ 59.6-59.6 $59.8$ $30.1$ 49.6-49.6 TAS-RL SR $59.1$ $55.2$ 6.6-6.6 $59.1$ $57.8$ -2.3 $59.1$ $55.2$ 6.7-6.7 $59.1$ $56.0$ 5.2-5.2 TA $41.6$ $37.3$ 10.3-10.3 $41.6$ $35.8$ 14.0-14.0 $41.6$ $32.1$ 22.9-22.9 $41.6$ $35.1$ 15.7-15.7 LLM [18] SR $28.5$ $43.3$ +34.1+34.1 $28.5$ $47.2$ +39.6+39.6 $28.5$ $34.2$ +16.7+16.7 $28.5$ $41.6$ +31.4+31.4 TA $24.3$ $30.7$ +20.9+20.9 $24.3$ $31.3$ +22.3+22.3 $24.3$ $21.7$ 10.8-10.8 $24.3$ $27.9$ +12.9+12.9 TAS-LLM SR $50.6$ $49.4$ -2.4 $50.6$ $53.2$ +4.8+4.8 $50.6$ $47.9$ -5.4 $50.6$ $50.2$ -0.9 TA $36.9$ $33.8$ -8.5 $36.9$ $34.6$ -6.3 $36.9$ $32.4$ 12.2-12.2 $36.9$ $33.6$ 9.0-9.0
5.3 Navigation Inference
Figure 4: Navigation Generalizability on DOMs (RCS): Relative Change in Success (RCS) [17] measures an agent’s adaptability, comparing dynamic to static TG performance (optimal RCS is $0\%$). Positive values indicate better dynamic performance (chance encounters), while negative values reflect poor adaptability. Both our TAS-enhanced agents approach optimal RCS, demonstrating better generalization to non-stationary environments.
Table 2 presents the results of our agents on the navigation task on DOMs. We run experiments on $10$ Matterport3D scans, each with $10$ randomly selected starting nodes $r\in G$, $20$ episodes, and each episode having $20$ timesteps. We draw $4$ key inferences from these observations to understand the generalizability of the various navigation approaches:
TAS improves Generalizability: Figure 4 presents a comparison of all approaches on RCS. TAS-LLM and TAS-RL both have a relatively low RCS score, meaning that navigation agents perform comparably in stationary graphs and DOMs. This highlights the effectiveness of incorporating transit-awareness in non-stationary targets. In contrast, the greedy heuristic Oracle agent and the zero-shot LLM lack transit information, with the former relying on privileged knowledge and latter utilizing commonsense cues about static observations for decision-making. We can infer the impact of TAS as a key component for the real-world deployment of embodied agents. Future work will explore the heuristic estimation of human habits as a prior for estimating transit paths for TAS.
[…]
Greedy Oracle Struggles With Shifting Objects: The Oracle agent uses a greedy heuristic to select a node towards the closest target object, and this works well in the static scenario. On DOMs however, targets are on the move and the greedy oracle tends to take a suboptimal path. The oracle heads in one particular direction towards a certain target object which was nearest, but keeps switching directions when another target appears to be closer. This behavior is a direct approximation of the distance reward with DD-PPO that causes target chasing and highlights the need for transit awareness.
Entropy Influences Performance: In all policies barring the Random agent baseline, we observe that the performance on a Fully Routine $\Lambda$ exhibiting low entropy is usually worse than the Semi-Routine or Random $\Lambda$ cases. This can be intuitively attributed to the agent encountering more targets in environments with higher object transit entropy (Fig. 3 in the Appendix). Despite this, we observe a strong performance of the TAS-RL agent in this case, with very similar values across all transit cases on SR and TA. This indicates that TAS can enhance embodied RL agents to become agnostic to transit scenarios where all target objects follow the same $\Lambda$. Future work will study the case of mixed $\Lambda$ with different objects following a different transit scenario.
6 Conclusion, Limitations, and Further Work
We present a novel approach to perform embodied navigation in dynamic environments with non-stationary targets. Our method enhances RL and LLM-based policies traditionally modeled for stationary targets with a Transit-Aware Strategy (TAS) that enables them to capture object transit information. To evaluate our TAS-enhanced agents, we introduce DOMs, or Dynamic Object Maps, an algorithm to translate static topological graphs into dynamic environments with shifting objects. Object transitions in DOMs aim to simulate human object placement habits. Our work breaks the underlying “static graph, stationary target” assumption common with embodied navigation literature, and presents a navigation benchmark in a dynamic environment with non-stationary targets. Our TAS-enhanced agents outperform non-TAS counterparts by $21.1\%$ when measured by average Success Rate in DOMs, while also drastically generalizing better to non-stationary environments by over $44.5\%$ when measured by Relative Change in Success. Code and data for our work will be made publicly available, and we hope to foster research in this domain. Future work will look at the sim2real transfer of TAS agents to real-world environments.
Supplementary Material
7 DOM Algorithm
To “DOMify” our environments, we use the following algorithm. Given a set of portable objects $\mathcal{O}$ and a topological graph $G$, we generate an evolving node-attributed graph trajectory $\zeta$ over $T$ timesteps that follows an object transit scenario $\Lambda$. Objects $\mathcal{O}$ and $\Lambda$ use Table 3 as reference. We manually populate this table with commonsense knowledge on general object locations. This grounded placement plays a key role later in influencing LLM agent performance, since they rely on LLM commonsense knowledge to make navigation decisions.
Algorithm 1 DOM: Modify $G$ With Object Transitions $\Lambda$
Input: Topological Graph $G=(V,E)$, Portable Object Set $\mathcal{O}$, Episode Length $T$, Initial Agent Node $r$, Object Transit Scenario $\Lambda$
Output: Evolving Graph Trajectory $\zeta$
Initialize: Add $\mathcal{O}$ portable objects to $G_{0}$.
1:for $t\in[1,\dots,T]$ do
2:  $G_{t}=G_{t-1}$
3:  for each object $o\in\mathcal{O}$ do
4:   Using Equation 1: $O(v,t)\leftarrow O(v,\Lambda)$
5:   Sample $v_{k}\in V$ from $O(v,t)$ {Following $\Lambda$}
6:   if $v_{k}\neq v_{o}$ then
7:    if $\Lambda=\text{Fully-Routine}$ then
8:     Use fixed trajectory $\delta(t)$
9:    else
10:     Compute random path $\delta(t)$ from $v_{o}$ to $v_{k}$
11:    end if
12:    Move $o$ along $\delta$ to $v_{k}$ over $t$ timesteps
13:   end if
14:  end for
15:  Add $G_{t}$ to $\zeta$
16:end for
17:return $\zeta$
Table 3: Rooms and Portable Objects: We map 2121 portable objects to a set of household rooms. This mapping is used to set destinations for object transit. During each episode, objects are placed in various rooms for a range of timesteps. Commonly moved objects such as phone, headphones are associated with $9$ different rooms, while less commonly shifted ones such as dumbbells appear only in the Gym.
Room Portable Objects
Bedroom Charger, Water Bottle, Smartwatch, Laptop, Notebook, Toothbrush, Mug, USB Flash Drive, Phone, Headphones, Hat
Garage Screwdriver, Flashlight, Mug, Phone, Headphones, Hat
Dining Salt and Pepper Shakers, Portable Speaker, Charger, Water Bottle, Mug, Bowl, Phone, Headphones, Hat
Office Charger, Laptop, Hat, Notebook, USB Flash Drive, Mug, Phone, Headphones
Bathroom Toothbrush, Phone, First-Aid Kit
Kitchen Salt and Pepper Shakers, Hat, Mug, Bowl, Phone, Headphones, First-Aid Kit
Lounge Playing Cards, Mug, Portable Speaker, Charger, Water Bottle, Laptop, Phone, USB Flash Drive, Dice, Headphones, Hat
Gym Dumbbells, Jumprope, Smartwatch, Phone, Headphones, Hat
Outdoor Jumprope, Smartwatch, Portable Speaker, Phone, Water Bottle, Headphones, Hat
Recreation Playing Cards, Dice, Water Bottle, Headphones, Hat
8 DOMifying MP3D
We convert MP3D graphs into DOMs with the $21$ different objects presented in Table 3. On each of the $10$ scans chosen, we introduce routine, semi-routine and random object movement, and provide these as dictionaries with varying seeds. We will release these modified environments along with supporting code.
Table 4: Summary statistics of scans used for Table 2, including each scan’s ID, number of nodes, edges, and rooms. If a scan has multiple rooms of the same type (e.g. two bathrooms), each instance is counted separately from the total.
Scan ID # of Nodes # of Edges # of Rooms
QUCTc6BB5sX 145 248 28
8194nk5LbLH 20 32 6
TbHJrupSAjP 114 221 28
2azQ1b91cZZ 215 531 30
oLBMNvg9in8 111 185 31
zsNo4HB9uLZ 53 84 17
EU6Fwq7SyZv 78 166 19
X7HyMhZNoso 84 143 25
x8F5xyUWy9e 43 86 10
Z6MFQCViBuw 58 91 18
8.1 Implementing Object Placement
In this section, we give more details on how we convert Matterport3D environments into DOMs.
Matterport3D Modifications:
Each Matterport3D (MP3D) scan represents a household environment consisting of a set of panoramic view points. Along with the viewpoints (or nodes), we are also given the exact 3D position as well as the relative distance between them. For modifying the MP3D environment, we first construct topological graphs of each scan, with nodes containing the position and the panoramic image, and edges containing relative distance between them. We consider 10 different scans chosen from the REVERIE [31] and R2R [1] unseen validation splits for inference. We choose scans according to these datasets as they contain a variety of rooms for us to populate (Refer Table 3 in the main paper).
The nodes of the topological graphs are then updated at each timestep with the portable objects according to the object placement scenario that has been chosen.
Strategy Overview:
We compute trajectories of the portable objects offline. For a fixed random seed $s$ and for each portable object $o_{p}\in O_{p}$, we create a sequence of nodes from the graph. In the random transit scenario, we choose any node in the graph. In the routine and semi-routine scenarios, we only choose nodes from plausible rooms. If the node chosen is not the same as the current node, we find the shortest path between the current node and the target node and add the nodes in the path to the sequence representing the trajectory. Once $o_{p}$ gets to the target node, it stays there for either $2$ or $3$ timesteps, determined by a random number generator. For each timestep $o_{p}$ stays at the node, we add the node to the trajectory sequence. After the staying period is over we select a new node and repeat the process until we hit $T$ timesteps.
We take the resulting trajectories and restructure the data to model an evolving graph over a period of $T$ timesteps. The resulting structure is a nested dictionary representing the changing graph. The keys of the outer dictionary are the timesteps [1,,T1,...,T]. The inner dictionary for each timestep $t$ has the node string ids as the keys and the values are the list of portable objects at that node at $T$. At each timestep $t\in[1,...,T]$ when the agent reaches a node, we simply use $t$ and the node id to retrieve the list of portable objects the agent is currently observing.
8.2 TAS-PPO
We modify a DD-PPO agent with temporal information about target objects in its observation space. As mentioned in section 3.3, we gather object paths $\mathcal{P}$ for augmenting this space.
In our task, since we treat each object equally, i.e. the objective is simply to maximize finding targets, we include $\mathcal{P}$ as a single list $l=[v_{1},v_{2},\dots,v_{n}]$ of size $n=|G|$. Each index of $l$ corresponds to a different node in $G$. For each index $i$ of $l$, $v_{i}$ is the number of unique target objects at the $i$-th node in $G$. $l$ is updated at each timestep, and passed to the observation space along with the agent’s current position and timestep. The agent’s action space is masked, with the entire graph being the action space, and the mask allowing only for selecting neighboring nodes.
Additionally, we swap out the distance reward $R_{dist}$ with the intersection reward $R_{I}$ described in the main text. $R_{I}$ here is triggered whenever the agent encounters a node where a target object has been in the last $5$ timesteps. We do not make any other modifications to the default setup.
8.3 TAS-LLM
We utilize the official implementation of LGX  [^1] https://github.com/vdorbala/LGX and incorporate it into our modified MP3D environment. Briefly, at each timestep, LGX scans the node for objects, and asks an LLM for directions to reach a target. In our case, we seek to maximize finding portable targets. As such, we prompt GPT-4 with the following base prompts -
System Prompt - “I am a smart robot trying to find as many portable objects as I can at home.”
User Prompt - “Which object from <<OBJECT_LIST>> should I go towards to find a new portable object? Reply in ONE word.”
The <<OBJECT_LIST>> here contains a set of objects that have been detected using YOLO-v8, as proposed in LGX. Additionally, we if a portable object is present at a certain node at a given timestep, we add it to this list.
The LLM then predicts a target object from the list, which is mapped to an adjacent node using our customized MP3D functions. The agent then moves to that node.
For the TAS-LGX case, we modify the System Prompt, with memory as presented in Section 3.3. We additionally ask -
System Prompt:
“I have seen the following objects and taken the following actions so far -
1. <<OBJECT_LIST>>: ACTION
2. <<OBJECT_LIST>>: ACTION
User Prompt:
“Which object from <<OBJECT_LIST>> should I go towards to find a new portable object? Reply in ONE word.”
Note that we maintain a horizon to avoid token overflow, where the earliest appended observation is removed when the context length limit is reached.
9 Video, Code and Dataset
We provide an anonymous link to our Code: https://anonymous.4open.science/r/otg-1AC0. The dataset will be present as a downloadable link in the code repository.
Furthermore, please see the attached video in the zip file for more information and demonstrations.